OpenStudy (anonymous):

Could someone please help me, the problem is 2 cube root 5 times 4 cube root 5 equals?

OpenStudy (anonymous):

Are you talking about \[2\sqrt[3]{5}*4\sqrt[3]{5}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay, then\[8\sqrt[3]{5^{2}}\] is your answer. Just multiply the 2 and 4 together. And add the exponents on the 5's

OpenStudy (anonymous):

ty that is what i had

OpenStudy (anonymous):

but i was worried 8 cube root 5 might be correct

OpenStudy (anonymous):

the 5 is squared its 5^(2/3) not just 5^(1/3)